a pipe open at both ends resonates to a frequency n1 and a pipe closed at one end resonates to a frequency n2. if they are joined to form a pipe closed at one end , then fundamental frequency will be .......
a) (n1n2)/(2n2+n1)
b) (2n2n1)/(2n2+n1)
c) (2n2n1)/n1+n2
d) (n2+2n1)/n1n2

Grade:12

1 Answers

Nirmal Singh.
10 years ago
For open pipe the; fundamental frequency is n1 = v / 2l1
l1` = v / 2n1
For closed pipe the fundamental frequency is n2 = v / 4l2
l2` = v / 4n2
For both pipe is joined and make a closed pipe length will be (l1 +l2) then fundamental frequency is n = v / 4(l1 + l2 )
by putting value of l1 and l2 then we will get n = n1n2 / (n1 + 2n2)
option a is correct
Thanks &Regards,
Nirmal Singh
Askiitians Faculty